Can you solve please by addition?
Thank you.
3x+4y=5
6x-4y=-14

Answer by uknouluvme89(30)

3x+4y=5
6x-4y=-14
multiply the top equation by -2
(-2)(3x+4y=5)
6x-4y=-14
the new system of equations will be:
-6x-8y=-10
+6x-4y=-14
add them together to get
-12y = -24
divide both sides by -12 to get y=2
substitute the value of y back into one of the original equations to find x
3x+4(2)=5
3x+8=5
3x=-3
x=-1
hope that was helpful!